# Communication-Efficient Stochastic Distributed Learning

### v1

@@ C-A Key bounds
Lemma 4.
Let Assumptions 1, 2 and 4 hold; when $\gamma\tau\leq\frac{2}{|\tilde{\lambda}_{\max}|\rho}$ and $6\gamma^{2}\tau(3L^{2}+\rho^{2}d_{u}^{2})<\frac{1/4}{\tau-1}$, we have
$\mathbb{E}[\|\widehat{\mathbf{\Phi}}_{k}\|^{2}]\leq\left(\frac{72\gamma\tau^{2}}{|\tilde{\lambda}_{\max}|\rho}+(18+\frac{36\gamma\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho})16\tau^{3}\gamma^{2}\right)\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]$ (49)
$+16\tau^{3}\gamma^{2}N\mathbb{E}[\|\nabla F(\bar{x}_{k})\|^{2}]+48\tau^{2}\gamma^{2}L^{2}\mathbb{E}[t_{k}]$
Proof.
@@
$\mathbb{E}[\left\|\Phi_{k}^{t+1}-\bar{\mathbf{X}}_{k}\right\|^{2}]$
$=\mathbb{E}[\|\Phi_{k}^{t}-\bar{\mathbf{X}}_{k}+\gamma\mathbf{Y}_{k}-\gamma(G(\Phi_{k}^{t})-\nabla\mathrm{F}(\bar{\mathbf{X}}_{k})+$
$+\rho\mathbf{D}\Phi_{k}^{t}-\rho\mathbf{D}\mathbf{X}_{k})\|^{2}]$
$\leq\left(1+\frac{1}{\tau-1}\right)\mathbb{E}[\left\|\Phi_{k}^{t}-\bar{\mathbf{X}}_{k}\right\|^{2}]+$
$+\tau\gamma^{2}\mathbb{E}[\|Y_{k}-(G(\Phi_{k}^{t})-\nabla\mathrm{F}(\bar{\mathbf{X}}_{k})+\rho\mathbf{D}\Phi_{k}^{t}-\rho\mathbf{D}\mathbf{X}_{k})\|^{2}]$
$\leq\left(1+\frac{1}{\tau-1}\right)\mathbb{E}[\left\|\Phi_{k}^{t}-\bar{\mathbf{X}}_{k}\right\|^{2}]+$
$+2\tau\gamma^{2}\mathbb{E}[\|G(\Phi_{k}^{t})-\nabla F(\Phi_{k}^{t})+\nabla F(\Phi_{k}^{t})-\nabla\mathrm{F}(\bar{\mathbf{X}}_{k})+$
$+\rho\mathbf{D}\Phi_{k}^{t}-\rho D\bar{\mathbf{X}}_{k}\|^{2}]+$
$+2\tau\gamma^{2}\mathbb{E}[\left\|\mathbf{Y}_{k}+\rho\mathbf{D}\bar{\mathbf{X}}_{k}-\rho\mathbf{D}\mathbf{X}_{k}\right\|^{2}]$
$\leq\left(1+\frac{1}{\tau-1}+6\gamma^{2}\tau(2L^{2}+L^{2}+\rho^{2}d_{u}^{2})\right)\mathbb{E}[\left\|\Phi_{k}^{t}-\bar{\mathbf{X}}_{k}\right\|^{2}]+$
$+2\tau\gamma^{2}\mathbb{E}[\left\|\mathbf{Y}_{k}+\rho\mathbf{D}\bar{\mathbf{X}}_{k}-\rho\mathbf{D}\mathbf{X}_{k}\right\|^{2}]+6\tau\gamma^{2}(2L^{2}\mathbb{E}[t^{t}_{k}])$
$\leq\left(1+\frac{5/4}{\tau-1}\right)\mathbb{E}[\|\Phi_{k}^{t}-\bar{\mathbf{X}}_{k}\|^{2}]+12\tau\gamma^{2}L^{2}\mathbb{E}[t^{t}_{k}]+$
$+2\tau\gamma^{2}\mathbb{E}[\|\mathbf{Y}_{k}+\rho\mathbf{D}\bar{\mathbf{X}}_{k}-\rho\mathbf{D}\mathbf{X}_{k}\|^{2}],$ (50)
@@
$6\gamma^{2}\tau(3L^{2}+\rho^{2}d_{u}^{2})\leq\frac{1/4}{\tau-1}.$ (51)
Iterating (50) for $t=0,...,\tau-1$ yields
$+2\tau\gamma^{2}\sum_{l=0}^{t}\left(1+\frac{5/4}{\tau-1}\right)^{l}\mathbb{E}[\|\mathbf{Y}_{k}+\rho\mathbf{D}\bar{\mathbf{X}}_{k}-\rho\mathbf{D}\mathbf{X}_{k}\|^{2}]+$
$+\sum_{l=0}^{t}\left(1+\frac{5/4}{\tau-1}\right)^{t-l}12\tau\gamma^{2}L^{2}\mathbb{E}[t_{k}^{l}]$
$\leq 4\mathbb{E}[\left\|\mathbf{X}_{k}-\bar{\mathbf{X}}_{k}\right\|^{2}]+8\tau^{2}\gamma^{2}\mathbb{E}[\left\|\mathbf{Y}_{k}+(\rho\mathbf{D}\bar{\mathbf{X}}_{k}-\rho\mathbf{D}\mathbf{X}_{k})\right\|^{2}]$
$+48\tau^{2}\gamma^{2}L^{2}\mathbb{E}[t^{t}_{k}],$ (52)
where in the last inequality we used $(1+\frac{a}{\tau-1})^{t}\leq\exp(\frac{at}{\tau-1})\leq\exp(a)$ for $t\leq\tau-1$. It is easy to verify that (52) also holds for $\tau=1$.
Summing (52) over $t$, and using (34), it follows that
$\mathbb{E}[\|\widehat{\mathbf{\Phi}}_{k}\|^{2}]\leq 4\tau\mathbb{E}[\|\mathbf{X}_{k}-\bar{\mathbf{X}}_{k}\|^{2}]+8\tau^{3}\gamma^{2}\mathbb{E}[\|\mathbf{Y}_{k}-\bar{\mathbf{Y}}_{k}+\bar{\mathbf{Y}}_{k}$
$+(\rho\mathbf{D}\bar{\mathbf{X}}_{k}-\rho\mathbf{D}\mathbf{X}_{k})\|^{2}]+48\tau^{2}\gamma^{2}L^{2}\mathbb{E}[t_{k}]$
$\leq 4\tau\mathbb{E}[\left\|\mathbf{X}_{k}-\bar{\mathbf{X}}_{k}\right\|^{2}]+16\tau^{3}\gamma^{2}\mathbb{E}[\left\|\bar{\mathbf{Y}}_{k}\right\|^{2}]$
$+16\tau^{3}\gamma^{2}\mathbb{E}[\left\|\mathbf{Y}_{k}-\bar{\mathbf{Y}}_{k}+(\rho D\bar{\mathbf{X}}_{k}-\rho D{\mathbf{X}}_{k})\right\|^{2}]$
$+48\tau^{2}\gamma^{2}L^{2}\mathbb{E}[t_{k}]$
$\leq 4\tau\mathbb{E}[\left\|\mathbf{X}_{k}-\bar{\mathbf{X}}_{k}\right\|^{2}]+16\tau^{3}\gamma^{2}N\mathbb{E}[\left\|\nabla F(\bar{x}_{k})\right\|^{2}]$
$+16\tau^{3}\gamma^{2}\mathbb{E}[\|\mathbf{Y}_{k}-\bar{\mathbf{Y}}_{k}+(\rho D\bar{\mathbf{X}}_{k}-\rho D{\mathbf{X}}_{k})\|^{2}]$
$+48\tau^{2}\gamma^{2}L^{2}\mathbb{E}[t_{k}]$
$\leq\left(\frac{72\gamma\tau^{2}}{|\tilde{\lambda}_{\max}|\rho}+(18+\frac{36\gamma\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho})16\tau^{3}\gamma^{2}\right)\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]$
$+16\tau^{3}\gamma^{2}N\mathbb{E}[\left\|\nabla F(\bar{x}_{k})\right\|^{2}]+48\tau^{2}\gamma^{2}L^{2}\mathbb{E}[t_{k}],$
which concludes the proof. ∎
@@
Lemma 5.
Let $\left\{t_{k}\right\}$ be the iterates generated by LT-ADMM-VR. If $\gamma\leq\min\{\bar{\gamma}_{8},\bar{\gamma}_{9},\bar{\gamma}_{10}\}$, we have for all $k\in\mathbb{N}$:
@@
$s_{0}\coloneqq\frac{36m_{u}\gamma\tau}{|\tilde{\lambda}_{\max}|\rho}+\frac{4\tau}{m_{l}}\left(18+\frac{36\gamma\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho}\right)(\gamma^{2}+\gamma)2m_{u}$
$s_{1}\coloneqq\left(\frac{72\gamma\tau^{2}}{|\tilde{\lambda}_{\max}|\rho}+(18+\frac{36\gamma\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho})16\tau^{3}\gamma^{2}\right)\frac{4m_{u}}{m_{l}}$ (54)
$s_{2}\coloneqq\frac{64m_{u}}{m_{l}}\tau^{3}\gamma^{2}N+\frac{8m_{u}}{m_{l}}(\gamma^{2}+\gamma)\tau N.$
Proof.
From Algorithm 1, $\forall k,r_{i,h,k}^{t+1}=r_{i,h,k}^{k}$ with probability $1-\frac{1}{m_{i}}$ and $r_{i,h,k}^{t+1}=\phi_{i,k}^{t+1}$ with probability $\frac{1}{m_{i}}$. Therefore, we have
$\mathbb{E}\left[t^{t+1}_{k}\right]=\frac{1}{m_{i}}\sum_{h=1}^{m_{i}}\mathbb{E}\left[\left\|\mathbf{r}_{h,k}^{t+1}-\bar{\mathbf{X}}_{k}\right\|^{2}\right]$
@@
$=\left(1-\frac{1}{m_{i}}\right)\frac{1}{m_{i}}\sum_{h=1}^{m_{i}}\mathbb{E}\left[\|\mathbf{r}_{h,k}^{t}-\bar{X}_{k}\|^{2}\right]$
$+\frac{1}{m_{i}}\mathbb{E}\left[\left\|\Phi_{k}^{t+1}-\bar{\mathbf{X}}_{k}\right\|^{2}\right].$
Now, denoting $q^{t}_{k}=\mathbf{Y}_{k}-(G(\Phi_{k}^{t})-\nabla\mathrm{F}(\bar{\mathbf{X}}_{k})+\rho\mathbf{D}\Phi_{k}^{t}-\rho\mathbf{D}\mathbf{X}_{k}),$ we have
$\left\|\Phi_{k}^{t+1}-\bar{X}_{k}\right\|^{2}=\left\|\Phi_{k}^{t+1}-\Phi_{k}^{t}+\Phi_{k}^{t}-\bar{X}_{k}\right\|^{2}$
$=\|\Phi_{k}^{t}-\bar{X}_{k}\|^{2}+2\gamma\langle\Phi_{k}^{t}-\bar{X}_{k},q^{t}_{k}\rangle+\gamma^{2}\|q^{t}_{k}\|^{2}$
$\leq(1+\gamma)\|\Phi_{k}^{t}-\bar{X}_{k}\|^{2}+(\gamma^{2}+\gamma)\|q^{t}_{k}\|^{2},$
and using (48) yields
$\mathbb{E}\left[\left\|q^{t}_{k}\right\|^{2}\right]$
$=\mathbb{E}[\|\mathbf{Y}_{k}-(\nabla F(\Phi_{k}^{t})-\nabla\mathrm{F}(\bar{\mathbf{X}}_{k})+\rho\mathbf{D}\Phi_{k}^{t}-\rho\mathbf{D}\mathbf{X}_{k})\|^{2}]$
$+\mathbb{E}\left[\|G(\Phi_{k}^{t})-\nabla F(\Phi_{k}^{t})\|^{2}\right]$
$\leq\mathbb{E}\left[\|G(\Phi_{k}^{t})-\nabla F(\Phi_{k}^{t})\|^{2}\right]$
$+2\mathbb{E}[\|\nabla F(\Phi_{k}^{t})-\nabla\mathrm{F}(\bar{\mathbf{X}}_{k})+\rho\mathbf{D}\Phi_{k}^{t}-\rho\mathbf{D}\bar{\mathbf{X}}_{k}\|^{2}]$
$+2\mathbb{E}[\left\|\mathbf{Y}_{k}-\bar{\mathbf{Y}}_{k}+\bar{\mathbf{Y}}_{k}+(\rho\mathbf{D}\bar{\mathbf{X}}_{k}-\rho\mathbf{D}\mathbf{X}_{k})\right\|^{2}]$
$\leq\mathbb{E}\left[\|G(\Phi_{k}^{t})-\nabla F(\Phi_{k}^{t})\|^{2}\right]+2(L^{2}+\rho^{2}d_{u}^{2})\mathbb{E}[\|\Phi_{k}^{t}-\bar{\mathbf{X}}_{k}\|^{2}$
$+4\left(18+\frac{36\gamma\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho}\right)\|\widehat{\mathbf{d}}_{k}\|^{2}+4N\|\nabla F(\bar{x}_{k})\|^{2}$
$\leq(4L^{2}+2\rho^{2}d_{u}^{2})\mathbb{E}[\|\Phi_{k}^{t}-\bar{\mathbf{X}}_{k}\|^{2}+4\left(18+\frac{36\gamma\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho}\right)\|\widehat{\mathbf{d}}_{k}\|^{2}$
$+2L^{2}\mathbb{E}[t^{t}_{k}]+4N\|\nabla F(\bar{x}_{k})\|^{2}.$
Therefore we have
$\mathbb{E}\left[t^{t+1}_{k}\mid\mathcal{F}^{k}\right]=\left(1-\frac{1}{m_{i}}\right)\frac{1}{m_{i}}\sum_{h=1}^{m_{i}}\|\mathbf{r}_{h,k}^{t}-\bar{X}_{k}\|^{2}$
$+\frac{1}{m_{i}}\|\Phi_{k}^{t+1}-\bar{X}_{k}\|^{2}$
$\leq(1-\frac{1}{m_{i}})t^{t}_{k}+\frac{1}{m_{i}}\left((1+\gamma)\|\Phi_{k}^{t}-\bar{X}_{k}\|^{2}+(\gamma^{2}+\gamma)\|q^{t}_{k}\|^{2}\right)$
$\leq\left(1-\frac{1}{m_{u}}+\frac{2L^{2}}{m_{l}}(\gamma^{2}+\gamma)\right)\mathbb{E}[t^{t}_{k}]+$
$+\left(\frac{1}{m_{l}}(1+\gamma)+\frac{4L^{2}+2\rho^{2}d_{u}^{2}}{m_{l}}(\gamma^{2}+\gamma)\right)\mathbb{E}[\|\Phi_{k}^{t}-\bar{\mathbf{X}}_{k}\|^{2}+$
$+\frac{4}{m_{l}}\left(18+\frac{36\gamma\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho}\right)(\gamma^{2}+\gamma)\|\widehat{\mathbf{d}}_{k}\|^{2}+$
$+\frac{4N}{m_{l}}(\gamma^{2}+\gamma)\|\nabla F(\bar{x}_{k})\|^{2}$
$\leq\left(1-\frac{1}{2m_{u}}\right)\mathbb{E}[t^{t}_{k}]+\frac{2}{m_{l}}\mathbb{E}[\|\Phi_{k}^{t}-\bar{\mathbf{X}}_{k}\|^{2}+$
$+\frac{4}{m_{l}}\left(18+\frac{36\gamma\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho}\right)(\gamma^{2}+\gamma)\|\widehat{\mathbf{d}}_{k}\|^{2}+$
$+\frac{4N}{m_{l}}(\gamma^{2}+\gamma)\|\nabla F(\bar{x}_{k})\|^{2}$ (55)
@@
$\frac{2L^{2}}{m_{l}}(\gamma^{2}+\gamma)<\frac{1}{2m_{u}},\quad\gamma+(4L^{2}+2\rho^{2}d_{u}^{2})(\gamma^{2}+\gamma)<1,$ (56)
which can be satisfied by choosing $\gamma<\min\{\bar{\gamma}_{8},\bar{\gamma}_{9}\}$. Iterating (55) for $t=0,...,\tau-1$ then yields:
$\mathbb{E}[t^{t}_{k}]\leq(1-\frac{1}{2m_{u}})^{t}\mathbb{E}[\left\|\mathbf{X}_{k}-\bar{\mathbf{X}}_{k}\right\|^{2}]$
$+\frac{4}{m_{l}}\left(18+\frac{36\gamma\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho}\right)(\gamma^{2}+\gamma)\sum_{l=0}^{t-1}(1-\frac{1}{2m_{u}})^{l}\|\widehat{\mathbf{d}}_{k}\|^{2}$
$+\frac{4N}{m_{l}}(\gamma^{2}+\gamma)\sum_{l=0}^{t-1}(1-\frac{1}{2m_{u}})^{l}\|\nabla F(\bar{x}_{k})\|^{2}$
$+\frac{2}{m_{l}}\sum_{l=0}^{t-1}(1-\frac{1}{2m_{u}})^{t-1-l}\mathbb{E}[\|\Phi_{k}^{l}-\bar{\mathbf{X}}_{k}\|^{2}$
$\leq 9\frac{2\gamma\tau}{|\tilde{\lambda}_{\max}|\rho}(1-\frac{1}{2m_{u}})^{t}\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]$
$+\frac{4N}{m_{l}}(\gamma^{2}+\gamma)\sum_{l=0}^{t-1}(1-\frac{1}{2m_{u}})^{l}\|\nabla F(\bar{x}_{k})\|^{2}$
$+\frac{2}{m_{l}}\sum_{l=0}^{t-1}(1-\frac{1}{2m_{u}})^{t-1-l}\mathbb{E}[\|\Phi_{k}^{l}-\bar{\mathbf{X}}_{k}\|^{2}$
$+\frac{4}{m_{l}}\left(18+\frac{36\gamma\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho}\right)(\gamma^{2}+\gamma)2m_{u}\|\widehat{\mathbf{d}}_{k}\|^{2}.$
@@
$\leq\left(\frac{36m_{u}\gamma\tau}{|\tilde{\lambda}_{\max}|\rho}+\frac{4\tau}{m_{l}}(18+\frac{36\gamma\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho})(\gamma^{2}+\gamma)2m_{u}\right)\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]$
$+\frac{4m_{u}}{m_{l}}\|\widehat{\mathbf{\Phi}}_{k}\|^{2}+\frac{4N}{m_{l}}(\gamma^{2}+\gamma)2m_{u}\tau\|\nabla F(\bar{x}_{k})\|^{2},$
@@
$\mathbb{E}[t_{k}]$ $\leq(s_{0}+s_{1})\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]+s_{2}\mathbb{E}[\|\nabla F(\bar{x}_{k})\|^{2}]+$
$+\frac{192m_{u}}{m_{l}}\tau^{2}\gamma^{2}L^{2}\mathbb{E}[t_{k}],$
where $s_{0}$, $s_{1}$ and $s_{2}$ are defined in (54).
Letting $\gamma<\bar{\gamma}_{10}$, it follows that
$\frac{192m_{u}}{m_{l}}\tau^{2}\gamma^{2}L^{2}<\frac{1}{2},$ (57)
and thus $\mathbb{E}[t_{k}]\leq 2(s_{0}+s_{1})\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]+2s_{2}\mathbb{E}[\|\nabla F(\bar{x}_{k})\|^{2}].$
Lemma 6.
Let Assumptions 1, 2 and 4 hold; when $\gamma<\min\{\bar{\gamma}_{1},\bar{\gamma}_{11}\}$, it holds that $\forall k\geq 0$
$\mathbb{E}[\|\widehat{\mathbf{d}}_{k+1}\|^{2}]\leq(\delta+\frac{\tilde{q}_{1}}{1-\delta})\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]$ (58)
$+\frac{\tilde{q}_{2}}{1-\delta}\mathbb{E}[\|\sum_{t}\overline{\nabla F}(\boldsymbol{\Phi}^{t}_{k})\|^{2}]+\frac{\tilde{q}_{3}}{1-\delta}\mathbb{E}[\|\nabla F(\bar{x}_{k})\|^{2}]+\tilde{q}_{4}\mathbb{E}[t_{k}],$

### v2

@@ 9.1 Key bounds
Lemma 9.8.
Let Assumptions 1 and 2 hold; when $\beta\tau\leq\frac{2}{\lambda_{u}\rho}$ and $\gamma\leq\bar{\gamma}_{7}$, we have
$\mathbb{E}[\|\widehat{\mathbf{\Phi}}_{k}\|^{2}]\leq\left(\frac{72\beta\tau^{2}}{\lambda_{l}\rho}+144\tau^{3}\beta^{2}\right)\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]$ (50)
[…]
Proof 9.9.
@@
$\mathbb{E}[\left\|\Phi_{k}^{t+1}-\bar{\mathbf{X}}_{k}\right\|^{2}]$ (51)
$=\mathbb{E}[\|\Phi_{k}^{t}-\bar{\mathbf{X}}_{k}+\beta\mathbf{Y}_{k}-\gamma(G(\mathrm{\Phi}_{k}^{t})-\nabla\mathrm{F}(\bar{\mathbf{X}}_{k}))\|^{2}]$
$\leq\left(1+\frac{1}{\tau-1}\right)\mathbb{E}[\left\|\Phi_{k}^{t}-\bar{\mathbf{X}}_{k}\right\|^{2}]$
$+\tau\mathbb{E}[\|\beta\mathbf{Y}_{k}-\gamma(G(\mathrm{\Phi}_{k}^{t})-\nabla\mathrm{F}(\bar{\mathbf{X}}_{k}))\|^{2}]$
$\leq\left(1+\frac{1}{\tau-1}+4\gamma^{2}\tau(2L^{2}+L^{2})\right)\mathbb{E}[\left\|\mathbf{\Phi}_{k}^{t}-\bar{\mathbf{X}}_{k}\right\|^{2}]$
$+2\tau\beta^{2}\mathbb{E}[\left\|\mathbf{Y}_{k}\right\|^{2}]+4\tau\gamma^{2}(2L^{2}\mathbb{E}[t^{t}_{k}])$
$\leq\left(1+\frac{5/4}{\tau-1}\right)\mathbb{E}[\|\mathbf{\Phi}_{k}^{t}-\bar{\mathbf{X}}_{k}\|^{2}]$
$+8\tau\gamma^{2}L^{2}\mathbb{E}[t^{t}_{k}]+2\tau\beta^{2}\mathbb{E}[\|\mathbf{Y}_{k}\|^{2}],$
@@
$4\gamma^{2}\tau(2L^{2}+L^{2})\leq\frac{1/4}{\tau-1},$ (52)
which can be satisfied when $\gamma\leq\bar{\gamma}_{7}$. Similar to Lemma 8.4, we can derive that (50) holds, which concludes the proof.
@@
Lemma 9.10.
Let $\left\{t_{k}\right\}$ be the iterates generated by LT-ADMM-VR. If $\beta\tau\leq\frac{2}{\lambda_{l}\rho}$ and $\gamma\leq\min\{\bar{\gamma}_{8},\bar{\gamma}_{9},\bar{\gamma}_{10}\}$, we have for all $k\in\mathbb{N}$:
@@
$s_{0}=\frac{36\beta\tau^{2}m_{u}}{\lambda_{l}\rho}+\frac{144\tau^{2}m_{u}}{m_{l}}\beta^{2}$
$s_{1}=\left(\frac{72\beta\tau^{2}}{\lambda_{l}\rho}+144\tau^{3}\beta^{2}\right)\frac{8m_{u}\tau}{m_{l}}$
$s_{2}=\frac{16N\gamma^{2}m_{u}\tau^{2}}{m_{l}}+\frac{8m_{u}\tau}{m_{l}}16\tau^{3}\gamma^{2}N.$ (54)
Proof 9.11.
From Algorithm 1, $\forall k,r_{i,h,k}^{t+1}=r_{i,h,k}^{k}$ with probability $1-\frac{1}{m_{i}}$ and $r_{i,h,k}^{t+1}=\phi_{i,k}^{t+1}$ with probability $\frac{1}{m_{i}}$, therefore,
$\mathbb{E}[t^{t+1}_{k}]=\frac{1}{m_{i}}\sum_{h=1}^{m_{i}}\mathbb{E}[\|\mathbf{r}_{h,k}^{t+1}-\bar{\mathbf{X}}_{k}\|^{2}]$
@@
$=\left(1-\frac{1}{m_{i}}\right)\frac{1}{m_{i}}\sum_{h=1}^{m_{i}}\mathbb{E}\left[\|\mathbf{r}_{h,k}^{t}-\bar{\mathbf{X}}_{k}\|^{2}\right]$
$+\frac{1}{m_{i}}\mathbb{E}[\|\Phi_{k}^{t+1}-\bar{\mathbf{X}}_{k}\|^{2}].$
Denote $q^{t}_{k}=\beta\mathbf{Y}_{k}-\gamma(G(\mathrm{\Phi}_{k}^{t})-\nabla\mathrm{F}(\bar{\mathbf{X}}_{k})),$ we have $\left\|\mathrm{\Phi}_{k}^{t+1}-\bar{\mathbf{X}}_{k}\right\|^{2}=\left\|\mathrm{\Phi}_{k}^{t+1}-\mathrm{\Phi}_{k}^{t}+\mathrm{\Phi}_{k}^{t}-\bar{\mathbf{X}}_{k}\right\|^{2}\leq 2\|\Phi_{k}^{t}-\bar{\mathbf{X}}_{k}\|^{2}+2\|q^{t}_{k}\|^{2},$ and
$\mathbb{E}[\|q^{t}_{k}\|^{2}]$
$\leq 2\gamma^{2}\mathbb{E}[\|G(\mathrm{\Phi}_{k}^{t})-\nabla\mathrm{F}(\bar{\mathbf{X}}_{k})\|^{2}]+2\beta^{2}\mathbb{E}[\left\|\mathbf{Y}_{k}\right\|^{2}]$
$\leq 4\gamma^{2}(2L^{2}+L^{2})\mathbb{E}[\left\|\mathrm{\Phi}_{k}^{t}-\bar{\mathbf{X}}_{k}\right\|^{2}]+2\beta^{2}\mathbb{E}[\left\|\mathbf{Y}_{k}\right\|^{2}]$
$+4\gamma^{2}(2L^{2}\mathbb{E}[t^{t}_{k}])$
$\leq 12\gamma^{2}L^{2}\left\|\mathrm{\Phi}_{k}^{t}-\bar{\mathbf{X}}_{k}\right\|^{2}+4\gamma^{2}N\|\nabla{F}(\bar{{x}}_{k})\|^{2}$
$+4\beta^{2}\|\mathbf{Y}_{k}-\mathbf{\bar{Y}}_{k}\|^{2}+8\gamma^{2}L^{2}\mathbb{E}[t^{t}_{k}],$
it follows that
$\mathbb{E}[t^{t+1}_{k}]=(1-\frac{1}{m_{i}})\frac{1}{m_{i}}\sum_{h=1}^{m_{i}}\|\mathbf{r}_{h,k}^{t}-\bar{\mathbf{X}}_{k}\|^{2}$ (55)
$+\frac{1}{m_{i}}\|\mathrm{\Phi}_{k}^{t+1}-\bar{\mathbf{X}}_{k}\|^{2}$
$\leq(1-\frac{1}{m_{i}})t^{t}_{k}+\frac{1}{m_{i}}\left(2\|\mathrm{\Phi}_{k}^{t}-\bar{\mathbf{X}}_{k}\|^{2}+2\|q^{t}_{k}\|^{2}\right)$
$\leq\left(1-\frac{1}{m_{u}}+\frac{16\gamma^{2}L^{2}}{m_{l}}\right)\mathbb{E}[t^{t}_{k}]$ (56)
$+\left(\frac{2}{m_{l}}+\frac{24\gamma^{2}L^{2}}{m_{l}}\right)\mathbb{E}[\|\mathrm{\Phi}_{k}^{t}-\bar{\mathbf{X}}_{k}\|^{2}$
$+\frac{72}{m_{l}}\beta^{2}\|\widehat{\mathbf{d}}_{k}\|^{2}+\frac{8N}{m_{l}}\gamma^{2}\|\nabla F(\bar{\mathbf{X}}_{k})\|^{2}$
$\leq\left(1-\frac{1}{2m_{u}}\right)\mathbb{E}[t^{t}_{k}]+\frac{4}{m_{l}}\mathbb{E}[\|\mathrm{\Phi}_{k}^{t}-\bar{\mathbf{X}}_{k}\|^{2}$ (57)
$+\frac{72}{m_{l}}\beta^{2}\|\widehat{\mathbf{d}}_{k}\|^{2}+\frac{8N}{m_{l}}\gamma^{2}\|\nabla F(\bar{\mathbf{X}}_{k})\|^{2}$ (58)
@@
$\frac{16\gamma^{2}L^{2}}{m_{l}}<\frac{1}{2m_{u}},\quad{24\gamma^{2}L^{2}}<2.$ (59)
Iterating (58) for $t=0,...,\tau-1$ then yields:
$\mathbb{E}[t^{t}_{k}]\leq(1-\frac{1}{2m_{u}})^{t}\mathbb{E}[\|\mathbf{X}_{k}-\bar{\mathbf{X}}_{k}\|^{2}]$
$+\frac{72}{m_{l}}\beta^{2}\sum_{l=0}^{t-1}(1-\frac{1}{2m_{u}})^{t-1-l}\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]$
$+\frac{8N\gamma^{2}}{m_{l}}\sum_{l=0}^{t-1}(1-\frac{1}{2m_{u}})^{l}\|\nabla F(\bar{x}_{k})\|^{2}$
$+\frac{4}{m_{l}}\sum_{l=0}^{t-1}(1-\frac{1}{2m_{u}})^{t-1-l}\mathbb{E}[\|\Phi_{k}^{l}-\bar{\mathbf{X}}_{k}\|^{2}$
$\leq\frac{36\beta\tau m_{u}}{\lambda_{l}\rho}\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]+\frac{16N\gamma^{2}m_{u}\tau}{m_{l}}\|\nabla F(\bar{x}_{k})\|^{2}$
$+\frac{8m_{u}\tau}{m_{l}}\sum_{l=0}^{t-1}\mathbb{E}[\|\mathrm{\Phi}_{k}^{l}-\bar{\mathbf{X}}_{k}\|^{2}+\frac{144m_{u}\tau}{m_{l}}\beta^{2}\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}].$
@@
$\leq\left(\frac{36\beta\tau^{2}m_{u}}{\lambda_{l}\rho}+\frac{144\tau^{2}m_{u}}{m_{l}}\beta^{2}\right)\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]+\frac{8m_{u}\tau}{m_{l}}\|\widehat{\mathbf{\Phi}}_{k}\|^{2}$
$+\frac{16N\gamma^{2}m_{u}\tau^{2}}{m_{l}}\|\nabla F(\bar{x}_{k})\|^{2},$
@@
$\mathbb{E}[t_{k}]$ $\leq(s_{0}+s_{1})\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]+s_{2}\mathbb{E}[\|\nabla F(\bar{x}_{k})\|^{2}]$
$+\frac{8m_{u}\tau}{m_{l}}32\tau^{2}\gamma^{2}L^{2}\mathbb{E}[t_{k}],$
where $s_{0}$, $s_{1}$ and $s_{2}$ are defined in (54). Letting
$\frac{8m_{u}\tau}{m_{l}}32\tau^{2}\gamma^{2}L^{2}<\frac{1}{2},$ (60)
and thus (53) holds. The conditions (59) and (60) hold when $\gamma\leq\min\{\bar{\gamma}_{8},\bar{\gamma}_{9},\bar{\gamma}_{10}\}$.
Lemma 9.12.
Let Assumptions 1 and 2 hold; when $\beta\tau\leq\frac{2}{\lambda_{u}\rho}$ and $\gamma<\min\{\bar{\gamma}_{1},\bar{\gamma}_{7}\}$, it holds that $\forall k\geq 0$
$\mathbb{E}[\|\widehat{\mathbf{d}}_{k+1}\|^{2}]\leq(\delta+\frac{\tilde{q}_{0}}{1-\delta})\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]$ (61)
$+\frac{\tilde{q}_{1}}{1-\delta}\mathbb{E}[\|\sum_{t}\overline{\nabla F}(\boldsymbol{\Phi}^{t}_{k})\|^{2}]+\frac{\tilde{q}_{2}}{1-\delta}\mathbb{E}[\|\nabla F(\bar{x}_{k})\|^{2}],$